On May 1, 2017, Pinkley Company sells office furniture for €80,000 cash. The office
furniture originally cost €200,000 when purchased on January 1, 2009. Depreciation is
recorded by the straight-line method over 10 years with a residual value of €20,000. What
depreciation expense should be recorded on this asset in 2017?
a. €6,000.
b. €6,667.
c. €9,000.
d. €18,000.

correct Ans. (A) . please show your work
Answer #1

Answer: Option ( A ) 6000

Explanation:

Depreciation per year = 200,000 - 20,000 / 10 = 18,000

2017 Depreciation = 18000 * 4/ 12 = 6000

As Furniture sold on may 1, 2017 Depreciation is calculated only for 4 months